# How Much Can We Forget about Data Contamination?

### v3

@@ Abstract
The leakage of benchmark data into the training data has emerged as a significant challenge for evaluating the capabilities of large language models (LLMs). In this work, we challenge the common assumption that small-scale contamination renders benchmark evaluations invalid. First, we experimentally quantify the magnitude of benchmark overfitting based on scaling along three dimensions: The number of model parameters (up to 1.6B), the number of times an example is seen (up to 144), and the number of training tokens (up to 40B). If model and data follow the Chinchilla scaling laws, minor contamination indeed leads to overfitting. At the same time, even 144 times of contamination can be forgotten if the training data is scaled beyond five times Chinchilla, a regime characteristic of many modern LLMs. Continual pre-training of OLMo-7B corroborates these results. Next, we study the impact of the weight decay parameter on example forgetting, showing that empirical forgetting occurs faster than the cumulative weight decay. This allows us to gauge the degree of forgetting in large-scale training runs, indicating that many LLMs, including Lllama 3 405B, have forgotten the data seen at the beginning of training.
@@ 1 Introduction
A core principle of machine learning is that a model should not be trained on the test set used for evaluation [21]. For foundation models trained on Internet-scale data, there are increasing concerns that this principle is violated due to the leakage of benchmark evaluation data into the training data [70, 57]. Indeed, many LLM developers have found overlap between their training data and the benchmark questions used for evaluation [10, 22]. What is more, research on memorization [12, 13] shows that text sequences from the training data are sometimes encoded within the model, including machine learning datasets [32, 47, 55, 8].
@@
While the fact that data contamination can lead to invalid performance evaluations is now well-established [52, 45, 71, 40], little is known about the precise conditions under which this is the case. The main reason for this is that the training data is usually unknown, and contamination identified via clever research designs that work around this restriction [27, 57, 20]. Because modern foundation models are sometimes trained for over a million gradient steps [22], it is unclear whether a single update on contaminated data at some point during training necessarily impacts downstream evaluations. And indeed, there is quite some evidence that language models need to see samples repeatedly to have any impact on the final model. For example, many papers on memorization have found that it occurs only when a sample is frequently repeated in the training data [14, 6, 37]. The same is true for research on knowledge acquisition, where a fact needs to be paraphrased many times before it is finally remembered by the model [2, 11, 15].
@@
In this work, we study the impact of data contamination in a controlled setting. This means we train language models from scratch on datasets where we explicitly insert contaminated examples [40]. We begin by quantifying how the overall magnitude of benchmark overfitting (or the cross-entropy loss of an observed sample) changes as we scale along three critical dimensions: (1) the number of model parameters, (2) the number of training tokens, and (3) the number of repetitions of an example in the training data (Section 4.1). Holding the other two dimensions fixed, we find that the effect of scaling is monotone in each dimension. First, similar to many other works, we find that the tendency of a model to overfit increases in the number of parameters [28, 73, 14]. Second, and this is also expected, we find a clear scaling in the number of repetitions, where more frequently repeated observations exhibit stronger overfitting [14, 37]. More surprisingly, we find that the effect of contamination can vanish as we increase the number of training tokens, up to the point where 12 repetitions of an entire dataset in the training data have no impact on the downstream evaluation on that same dataset.
@@
Our investigation reveals that the forgetting dynamics of gradient descent [66, 39] is the reason why increasing the number of tokens alleviates the impact of contamination. Concretely, we show that training on five times Chinchilla [35] of clean data can cause a model to forget even 144 times repeated training examples (Section 4.2). Forgetting also occurs for larger models, a point that we demonstrate using OLMo-7B [30] (Section 4.3). What is the reason for the forgetting? We show that exposure to novel data is important. Interestingly, models tend to exhibit the strongest overfitting on examples seen repeatedly throughout training, even compared to those seen during the end (Section 4.2).
@@
Because running pre-training experiments is expensive, we also ask to what degree forgetting can be explained by the training dynamics of gradient descent. We find that the weight decay parameter and learning rate schedule of the AdamW optimizer [49] play a key part in forgetting past training examples (Section 5). Concretely, we demonstrate that the cumulative weight decay (Section 5.1) bounds empirical forgetting (Section 5.2). This approach allows us to gauge the degree of forgetting in large-scale training runs by analyzing the optimization hyperparameters (Section 5.3). It also allows us to approximate how the final model parameters balance the gradient updates from different stages of training. Our analysis indicates that many LLMs, including OLMo-7B and Llama 3 405B [22], have forgotten the data seen at the beginning of training.
@@
Taken together, our main contribution is to show that the impact of individual examples in the training data depends on the precise characteristics of the setting. There are settings where the effect can be significant; Chinchilla training is an important example (Section 4.1). However, there are equally realistic settings where individual examples dont’t matter - including quite likely the data-intensive training runs of many recent LLMs [25].
@@ 2 Related Work
Data Contamination. The GPT-3 paper [10] uses an n-gram-based approach to differentiate between “clean” and “dirty” benchmark questions. This approach has since been used in many LLM reports [17, 68], including Llama 3 [22], where it is estimated that there might be a performance gain of up to 8 and 14 percentage points on PiQA and HellaSwag, respectively. The GPT-4 technical report [1] remarkably concluded that “contamination overall has very little effect on the reported results”. This has since given rise to a literature that aims to detect [57], mitigate [46], and estimate the effect of [71, 8] data contamination under various assumptions, but crucially without access to the training data. This literature often challenges the conclusion that contamination overall has little effect in GPT-4 [70].
Forgetting. In machine learning, the term forgetting is frequently associated with “catastrophic” forgetting, where learning new tasks hurt the performance at previously solved tasks [48]. In the context of LLMs, catastrophic forgetting can occur during fine-tuning [51] or continual learning [36]. In contrast, this paper studies forgetting as a potential “natural” phenomenon of learning [67]. Tirumala et al. [66] study forgetting in language modeling and find, similar to Toneva et al. [67], that forgetting can be exponentially slow. In contrast, Jagielski et al. [39] find that models empirically do forget examples over time. Pagliardini et al. [58] propose to add a second momentum term to the AdamW optimizer, and show that this slows down the forgetting of past gradients.
Data Attribution. Data attribution methods [42, 38, 59] aim to identify data points responsible for specific model behaviors. We ask how much a model’s benchmark performance is influenced by seeing the example during training, which broadly falls within this field [31, 16]. Importantly, we directly measure the influence of contaminated examples through explicit retraining, avoiding the approximation errors that can occur when using data attribution methods [42, 26] for large-scale models [5, 4]
@@ 3 Background and Methods
While it is well known that an individual data point can be influential if the training data and model are small [42], we are concerned with the large-data regime where the influence of any individual example may vanish [5, 4]. To further clarify this setup, Section 3.1 gives a brief overview of recent developments in scaling the training data of LLMs. Section 3.2 details the used benchmarks and explains how we contaminate the training data. Section 3.3 discusses the problem of near-duplicate benchmark questions.
Models and Training Data. We train language models of up to 1.6B parameters using the architecture and hyperparameters from the GPT-3 paper [10]. For this, we adopt the llm.c codebase. The training data is the 100BT split of the FineWeb-Edu dataset [50]. We also train OLMo-7B [30] using the corresponding code and data [65].
@@ 3.1 We Consider the Regime of n-times Chinchilla
According to the Chinchilla scaling law, for every doubling of model size, the number of training tokens should also be doubled. The Chinchilla model itself has 70 billion (B) parameters and was trained on 1.4 trillion (T) tokens; suggesting that the number of training tokens should be roughly 20x the number of model parameters [35]. While the Chinchilla paper was highly influential, modern language models are trained on significantly more tokens [64]. For example, the OLMo-7B model was trained on 2.46T tokens, 17.5x the amount suggested by Chinchilla [30]. Similarly, the Llama 3 70B model was reportedly trained on 15T tokens, at over 10x Chinchilla [22, 54]. The same holds for almost all recent LLMs at the 7B parameter scale [25]. In this paper, we count the number of tokens a model is trained on as a multiple of its Chinchilla tokens.
3.2 We Evaluate on a Mix of Different Benchmarks
We evaluate the impact of data contamination using a mix of seven different benchmarks: ARC-Easy [19], Social-I-QA [63], WinoGrande [62], PiQA [7], BoolQ [18], MMLU [34], and HellaSwag [72]. This means that every evaluation contains questions from all seven benchmarks. To construct the mixed contamination data, we first concatenate the different benchmarks. We then partition the set of all benchmark questions into subsets ranging from 10,000 to 2,000 questions so that each subset contains all benchmarks in equal weight: HellaSwag: 19.58%, SocialIQA: 8.27%, PiQA: 19.7%, MMLU: 21.82%, BoolQ: 6.48%, ARC-Easy: 5.92%, and WinoGrande: 18.16%. A holdout set of 10,000 benchmark questions is never added to the training data. The other subsets are added to the training data, repeated either 4, 12, 36, or 144 times. [^1] In preliminary experiments, we found that these numbers pleasantly cover the range from statistically significant contamination to complete overfitting. We also considered repeating observations a single time, as in [40]. However, we found this often leads to accuracy differences of about one or two percentage points, just within the margin of our confidence intervals, which is undesirable. We consider exact contamination, that is we contaminate the training data with the same texts that the model is later evaluated on. We insert benchmark questions individually and at random positions into the training data. Models are evaluated zero-shot via the likelihood assigned to different sentence completions [23]. For more discussion and details about how contamination is performed, see Supplement A.1 and Supplement A.2.
3.3 We Filter Near-Duplicate Benchmark Questions
Our method requires no side effects from contaminating the training data with one question on the evaluation of another question. However, upon closer inspection, it turns out that all the commonly used benchmarks from the literature contain questions that are either near-duplicates or where the context of one question contains the answer to another question (for example, because the same text document was used to create multiple questions). To address this problem, we perform extensive filtering, removing duplicate questions where the length-normalized Levenshtein distance falls below a certain threshold [43, 56]. The issue of duplicate benchmark questions is detailed in Supplement A.3, where we also describe an experiment to verify that near-duplicate questions do not invalidate our method.
4 Experimental Results
We now present our main experimental results. We being in Section 4.1 by discussing the scaling in model parameters, training tokens, and repetitions in the training data. The following Section 4.2 discusses various experiments on forgetting. The first two sections rely on training small GPT-3 models. Section 4.3 complements this with an analysis of OLMo-7B [30].
4.1 Contamination Scales with Model, Data, and Repetitions
We conduct three different experiments to understand how the effect of data contamination scales with the number of model parameters, training tokens, and the number of times a contaminated example is seen. First, we train increasingly large models on the same dataset of 7B tokens. Second, we train 124M parameter models on increasingly many tokens. Third, we train increasingly large models according to the Chinchilla scaling laws [35], meaning that the number of training tokens scales linearly with the model parameters. In all experiments, we contaminate the training data uniformly at random with benchmark questions.
Figure 1 depicts the results of all three experiments. Because we are interested in the performance difference between the holdout data and the contaminated examples, Figure 1 depicts the accuracy gap between the holdout and contaminated examples in percentage points. In Figure 1(a), we see that the accuracy gap due to contamination is increasing in the number of model parameters. For a 124M parameter model trained on 7B tokens, the overfitting due to 4 times contamination is 5 percentage points. For a 1.6B parameter model train on the same dataset, it is 20. Next, Figure 1(b) shows that the accuracy gap is decreasing in the number of training tokens. For a 124M parameter model trained at 2x Chinchilla, the accuracy gap due to 12 times contamination is 18 percentage points. For a 124M parameter model trained at 15x Chinchilla, the same accuracy gap is within the confidence interval of the holdout. From Figure 1, we also see that the accuracy gap is increasing in the number of times an example is repeated. For a 350M parameter model trained on 7B tokens, the accuracy gap is 11, 25, 44, and 51 percentage points for 4, 12, 32, and 144 times repeated contamination, respectively.
@@ 4.1 Contamination Scales with Model, Data, and Repetitions
Because the accuracy gap increases in the number of model parameters and decreases in the number of tokens, the interesting question is how it behaves if model parameters and tokens are scaled jointly. A natural starting point is to double the number of training tokens for every doubling of model parameters, as specified by the Chinchilla scaling laws [35]. Figure 1(c) depicts the accuracy gap due to contamination as we train increasingly large Chinchilla-optimal models. While there is no clear monotone pattern, we see that moderate amounts of contamination can lead to significant overfitting. For the 774M parameter model, 4 times repeated contamination leads to an accuracy gap of 15 percentage points, suggesting that under Chinchilla training, a single time of contamination can lead to overfitting of as much as 3 percentage points. [^2] To contaminate the training data of a 774M model a single time with 10,000 benchmark questions, we need to insert $\sim$0.5 million tokens into 15.5 billion training tokens, about 0.003% of the training data.
4.2 Contamination Can be Completely Forgotten
@@
(b) After 3 Epochs
(c) After 7 Epochs
@@
Figure 2: The forgetting dynamic of neural network training. (a) The development of the cross-entropy loss difference between contaminated and holdout benchmark questions over the course of training. Contamination occurs between the first and second Chinchilla (1 and 2 on the x-axis). (b) Accuracy gaps after training for 3 Chinchilla. (c) Accuracy gaps after training for 7 Chinchilla. (d) Same as (a). (e)+(f) The accuracy gap depends on the average position of an example in the training data. Mean and bootstrapped 90% confidence intervals.
In the previous Section 4.1, we saw that the accuracy gap due to contamination decreases in the number of tokens up to the point where even 12 repetitions of a benchmark question in the training data can become insignificant. In this Section, we identify the forgetting dynamic of neural network training as the reason for this effect. We discuss how quickly forgetting occurs, whether examples are completely forgotten, and what kind of repetition makes a model remember.

### v4

@@ Abstract
The leakage of benchmark data into the training data has emerged as a significant challenge for evaluating the capabilities of large language models (LLMs). In this work, we challenge the common assumption that small-scale contamination renders benchmark evaluations invalid. First, we experimentally quantify the magnitude of benchmark overfitting based on scaling along three dimensions: The number of model parameters (up to 1.6B), the number of times an example is seen (up to 144), and the number of training tokens (up to 40B). If model and data follow the Chinchilla scaling laws, minor contamination indeed leads to overfitting. At the same time, even 144 times of contamination can be forgotten if the training data is scaled beyond five times Chinchilla, a regime characteristic of many modern LLMs. Continual pre-training of OLMo-7B corroborates these results. Next, we study the impact of the weight decay parameter on example forgetting, showing that empirical forgetting occurs faster than the cumulative weight decay. This allows us to gauge the degree of example forgetting in large-scale training runs, indicating that many LLMs, including Llama 3 405B, have forgotten the data seen at the beginning of training.
Machine Learning, ICML
@@ 1 Introduction
A core principle of machine learning is that a model should not be trained on the test set used for evaluation (Donoho, 2017). For foundation models trained on Internet-scale data, there are increasing concerns that this principle is violated due to the leakage of benchmark evaluation data into the training data (Xu et al., 2024; Oren et al., 2024). Indeed, many LLM developers have found overlap between their training data and the benchmark questions used for evaluation (Brown et al., 2020; Dubey et al., 2024). What is more, research on memorization (Carlini et al., 2019, 2021) shows that text sequences from the training data are sometimes encoded within the model, including machine learning datasets (Grynbaum & Mac, 2023; Liang et al., 2023; Nasr et al., 2023; Bordt et al., 2024).
@@
While the fact that data contamination can lead to invalid performance evaluations is now well-established (Magar & Schwartz, 2022; Li & Flanigan, 2024; Yang et al., 2023; Jiang et al., 2024), little is known about the precise conditions under which this is the case. The main reason for this is that the training data is usually unknown, and contamination identified via clever research designs that work around this restriction (Golchin & Surdeanu, 2024; Oren et al., 2024; Deng et al., 2024). Because modern foundation models are sometimes trained for over a million gradient steps (Dubey et al., 2024), it is unclear whether a single update on contaminated data at some point during training necessarily impacts downstream evaluations. And indeed, there is quite some evidence that language models need to see samples repeatedly to have any impact on the final model. For example, many papers on memorization have found that it occurs only when a sample is frequently repeated in the training data (Carlini et al., 2023; Biderman et al., 2023; Huang et al., 2024b). The same is true for research on knowledge acquisition, where a fact needs to be paraphrased many times before it is finally remembered by the model (Allen-Zhu & Li, 2023; Cao et al., 2024; Chang et al., 2024).
@@
In this work, we study the impact of data contamination in a controlled setting. [^1] Code is available at https://github.com/tml-tuebingen/forgetting-contamination/. This means we train language models from scratch on datasets where we explicitly insert contaminated examples (Jiang et al., 2024). We begin by quantifying how the overall magnitude of benchmark overfitting (or the cross-entropy loss of an observed sample) changes as we scale along three critical dimensions: (1) the number of model parameters, (2) the number of training tokens, and (3) the number of repetitions of an example in the training data (Section 4.1). Holding the other two dimensions fixed, we find that the effect of scaling is monotone in each dimension. First, similar to many other works, we find that the tendency of a model to overfit increases in the number of parameters (Goodfellow et al., 2016; Zhang et al., 2017; Carlini et al., 2023). Second, and this is also expected, we find a clear scaling in the number of repetitions, where more frequently repeated observations exhibit stronger overfitting (Carlini et al., 2023; Huang et al., 2024b). More surprisingly, we find that the effect of contamination can vanish as we increase the number of training tokens, up to the point where 12 repetitions of an entire dataset in the training data have no impact on the downstream evaluation on that same dataset.
@@
Our investigation reveals that the forgetting dynamics of gradient descent (Tirumala et al., 2022; Jagielski et al., 2023) is the reason why increasing the number of tokens alleviates the impact of contamination. Concretely, we show that training on five times Chinchilla (Hoffmann et al., 2022) of clean data can cause a model to forget even 144 times repeated training examples (Section 4.2). Forgetting also occurs for larger models, a point that we demonstrate using OLMo-7B (Groeneveld et al., 2024) (Section 4.3). What is the reason for the forgetting? We show that exposure to novel data is important. Interestingly, models tend to exhibit the strongest overfitting on examples seen repeatedly throughout training, even compared to those seen during the end (Section 4.2).
@@
Because running pre-training experiments is expensive, we also ask to what degree forgetting can be explained by the training dynamics of gradient descent. We find that the weight decay parameter and learning rate schedule of the AdamW optimizer (Loshchilov & Hutter, 2019) play a key part in forgetting past training examples (Section 5). Concretely, we demonstrate that the cumulative weight decay (Section 5.1) bounds empirical forgetting (Section 5.2). This approach allows us to gauge the degree of forgetting in large-scale training runs by analyzing the optimization hyperparameters (Section 5.3). It also allows us to approximate how the final model parameters balance the gradient updates from different stages of training. Our analysis indicates that many LLMs, including OLMo-7B and Llama 3 405B (Dubey et al., 2024), have forgotten the data seen at the beginning of training.
@@
Taken together, our main contribution is to show that the impact of individual examples in the training data depends on the precise characteristics of the setting. There are settings where the effect can be significant; Chinchilla training is an important example (Section 4.1). However, there are equally realistic settings where individual examples don’t matter - including quite likely the data-intensive training runs of many recent LLMs (Gemma Team, 2024).
Supplement Sections A.1 and A.2 provide an overview of the most important takeaways and limitations of our work.
@@ 2 Related Work
Data Contamination. The GPT-3 paper (Brown et al., 2020) uses an n-gram-based approach to differentiate between “clean” and “dirty” benchmark questions. This approach has since been used in many LLM reports (Chowdhery et al., 2023; Touvron et al., 2023), including Llama 3 (Dubey et al., 2024), where it is estimated that there might be a performance gain of up to 8 and 14 percentage points on PiQA and HellaSwag, respectively. The GPT-4 technical report (Achiam et al., 2023) remarkably concluded that “contamination overall has very little effect on the reported results”. This has since given rise to a literature that aims to detect (Oren et al., 2024), mitigate (Li et al., 2024), and estimate the effect of (Yang et al., 2023; Bordt et al., 2024) data contamination under various assumptions, but crucially without access to the training data. This literature often challenges the conclusion that contamination overall has little effect in GPT-4 (Xu et al., 2024). In closely related work, Jiang et al. (2024) show that inserting benchmark questions into the pre-training data can lead to overfitting. Kocyigit et al. (2025) investigate the impact of data contamination on machine translation.
Forgetting. In machine learning, the term forgetting is frequently associated with “catastrophic” forgetting, where learning new tasks hurt the performance at previously solved tasks (Lopez-Paz & Ranzato, 2017). In the context of LLMs, catastrophic forgetting can occur during fine-tuning (Luo et al., 2023) or continual learning (Huang et al., 2024a). In contrast, this paper studies forgetting as a potential “natural” phenomenon of learning (Toneva et al., 2019). Tirumala et al. (2022) study forgetting in language modeling and find, similar to Toneva et al. (2019), that forgetting can be exponentially slow. In contrast, Jagielski et al. (2023) find that models empirically do forget examples over time. Pagliardini et al. (2025) propose to add a second momentum term to the AdamW optimizer, and show that this slows down the forgetting of past gradients.
Data Attribution. Data attribution methods (Koh & Liang, 2017; Ilyas et al., 2022; Park et al., 2023) aim to identify data points responsible for specific model behaviors. We ask how much a model’s benchmark performance is influenced by seeing the example during training, which broadly falls within this field (Grosse et al., 2023; Choe et al., 2024). Importantly, we directly measure the influence of contaminated examples through explicit retraining, avoiding the approximation errors that can occur when using data attribution methods (Koh & Liang, 2017; Ghorbani & Zou, 2019) for large-scale models (Basu et al., 2021; Bae et al., 2022)
@@ 3 Background and Methods
While it is well known that an individual data point can be influential if the training data and model are small (Koh & Liang, 2017), we are concerned with the large-data regime where the influence of any individual example may vanish (Basu et al., 2021; Bae et al., 2022). To further clarify this setup, Section 3.1 gives a brief overview of recent developments in scaling the training data of LLMs. Section 3.2 details the used benchmarks and explains how we contaminate the training data. Section 3.3 discusses the problem of near-duplicate benchmark questions.
Models and Training Data. We train language models of up to 1.6B parameters using the architecture and hyperparameters from the GPT-3 paper (Brown et al., 2020, Table 2.1). For this, we adopt the llm.c codebase. The training data is the 100BT split of the FineWeb-Edu dataset (Lozhkov et al., 2024). We also train OLMo-7B (Groeneveld et al., 2024) using the corresponding code and data (Soldaini et al., 2024).
@@ 3.1 We Consider the Regime of n-times Chinchilla
According to the Chinchilla scaling law, for every doubling of model size, the number of training tokens should also be doubled. The Chinchilla model itself has 70 billion (B) parameters and was trained on 1.4 trillion (T) tokens; suggesting that the number of training tokens should be roughly 20x the number of model parameters (Hoffmann et al., 2022, Table 3). While the Chinchilla paper was highly influential, modern language models are trained on significantly more tokens (Sardana & Frankle, 2024). For example, the OLMo-7B model was trained on 2.46T tokens, 17.5x the amount suggested by Chinchilla (Groeneveld et al., 2024). Similarly, the Llama 3 70B model was reportedly trained on 15T tokens, at over 10x Chinchilla (Dubey et al., 2024; Meta AI, 2024). The same holds for almost all recent 7B-parameter LLMs (Gemma Team, 2024). In this paper, we count the number of tokens a model is trained on as a multiple of its Chinchilla tokens.
@@
3.2 We Evaluate on a Mix of Different Benchmarks
We evaluate the impact of data contamination using a mix of seven different benchmarks: ARC-Easy (Clark et al., 2018), Social-I-QA (Sap et al., 2019), WinoGrande (Sakaguchi et al., 2021), PiQA (Bisk et al., 2020), BoolQ (Clark et al., 2019), MMLU (Hendrycks et al., 2021), and HellaSwag (Zellers et al., 2019). This means that every evaluation contains questions from all seven benchmarks. To construct the mixed contamination data, we first concatenate the different benchmarks. We then partition the set of all benchmark questions into subsets ranging from 10,000 to 2,000 questions so that each subset contains all benchmarks in equal weight: HellaSwag: 19.58%, SocialIQA: 8.27%, PiQA: 19.7%, MMLU: 21.82%, BoolQ: 6.48%, ARC-Easy: 5.92%, and WinoGrande: 18.16%. A holdout set of 10,000 benchmark questions is never added to the training data. The other subsets are added to the training data, repeated either 4, 12, 36, or 144 times. [^2] In preliminary experiments, we found that these numbers pleasantly cover the range from statistically significant contamination to complete overfitting. We also considered repeating observations a single time, as in (Jiang et al., 2024). However, we found this often leads to accuracy differences of about one or two percentage points, just within the margin of our confidence intervals, which is undesirable. We consider exact contamination, that is we contaminate the training data with the same texts that the model is later evaluated on. We insert benchmark questions individually and at random positions into the training data. Models are evaluated zero-shot via the likelihood assigned to different sentence completions (Gao, 2021). For more discussion and details about how contamination is performed, see Supplement A.3 and Supplement A.4.
3.3 We Filter Near-Duplicate Benchmark Questions
Our method requires no side effects from contaminating the training data with one question on the evaluation of another question. However, upon closer inspection, it turns out that all the commonly used benchmarks from the literature contain questions that are either near-duplicates or where the context of one question contains the answer to another question (for example, because the same text document was used to create multiple questions). To address this problem, we perform extensive filtering, removing duplicate questions where the length-normalized Levenshtein distance falls below a certain threshold (Levenshtein, 1966; Navarro, 2001). The issue of duplicate benchmark questions is detailed in Supplement A.5, where we also describe an experiment to verify that near-duplicate questions do not invalidate our method.
@@ 3.3 We Filter Near-Duplicate Benchmark Questions
(b) After 3 Chinchilla
(c) After 7 Chinchilla
@@
Figure 2: The forgetting dynamic of neural network training. (a) The development of the cross-entropy loss difference between contaminated and holdout benchmark questions over the course of training. Contamination occurs between the first and second Chinchilla (1 and 2 on the x-axis). (b) Accuracy differences after training for 3 Chinchilla. (c) Accuracy differences after training for 7 Chinchilla. (d) Same as (a). (e)+(f) The accuracy difference depends on the average position of an example in the training data. Mean and bootstrapped 90% confidence intervals.
4 Experimental Results
We now present our main experimental results. We begin in Section 4.1 by discussing the scaling in model parameters, training tokens, and repetitions in the training data. The following Section 4.2 discusses various experiments on forgetting. The first two sections rely on training small GPT-3 models. Section 4.3 complements this with an analysis of OLMo-7B (Groeneveld et al., 2024).
4.1 Contamination Scales with Model, Data, and Repetitions
We conduct three different experiments to understand how the effect of data contamination scales with the number of model parameters, training tokens, and the number of times a contaminated example is seen. First, we train increasingly large models on the same dataset of 7B tokens. Second, we train 124M parameter models on increasingly many tokens. Third, we train increasingly large models according to the Chinchilla scaling laws (Hoffmann et al., 2022), meaning that the number of training tokens scales linearly with the model parameters. In all experiments, we contaminate the training data uniformly at random with benchmark questions.
Figure 1 depicts the results of all three experiments. Because we are interested in the performance difference between the holdout data and the contaminated examples, Figure 1 depicts the accuracy difference between the holdout and contaminated examples in percentage points. In Figure 1(a), we see that the accuracy difference due to contamination is increasing in the number of model parameters. For a 124M parameter model trained on 7B tokens, the overfitting due to 4 times contamination is 5 percentage points. For a 1.6B parameter model train on the same dataset, it is 20. Next, Figure 1(b) shows that the accuracy difference is decreasing in the number of training tokens. For a 124M parameter model trained at 2x Chinchilla, the accuracy difference due to 12 times contamination is 18 percentage points. For a 124M parameter model trained at 15x Chinchilla, the respective accuracy difference is within the confidence interval of the holdout. From Figure 1, we also see that the accuracy difference is increasing in the number of times an example is repeated. For a 350M parameter model trained on 7B tokens, the accuracy difference is 11, 25, 44, and 51 percentage points for 4, 12, 32, and 144 times repeated contamination, respectively.
Because the accuracy difference increases in the number of model parameters and decreases in the number of tokens, the interesting question is how it behaves if model parameters and tokens are scaled jointly. A natural starting point is to double the number of training tokens for every doubling of model parameters, as specified by the Chinchilla scaling laws (Hoffmann et al., 2022). Figure 1(c) depicts the accuracy difference due to contamination as we train increasingly large Chinchilla-optimal models. While there is no clear monotone pattern, we see that moderate amounts of contamination can lead to significant overfitting. For the 774M parameter model, 4 times repeated contamination leads to an accuracy difference of 15 percentage points, suggesting that under Chinchilla training, a single time of contamination can lead to overfitting of as much as 3 percentage points. [^3] To contaminate the training data of a 774M model a single time with 10,000 benchmark questions, we need to insert $\sim$0.5 million tokens into 15.5 billion training tokens, about 0.003% of the training data.
4.2 Contamination Can be Completely Forgotten
In the previous Section 4.1, we saw that the accuracy difference due to contamination decreases in the number of tokens up to the point where even 12 repetitions of a benchmark question in the training data can become insignificant. In this Section, we identify the forgetting dynamic of neural network training as the reason for this effect. We discuss how quickly forgetting occurs, whether examples are completely forgotten, and what kind of repetition makes a model remember.